OpenStudy (anonymous):

Relationship A has a greater rate than Relationship B. The table represents Relationship B. Which equation could represent Relationship A? Choose ALL answers that are correct. Hours worked 2 4 5 8 Amount paid ($) 25 50 62.50 100 A. Y= 11.5x B. Y=12.25x C. Y=12.75x Y=13x @iGreen

OpenStudy (anonymous):

Can you help me with this one @iGreen?

OpenStudy (texaschic101):

let me get this clear....Is the number under the 4 a 62 and the number under the 5 a 50 ?

OpenStudy (texaschic101):

or is that a decimal in there ??

OpenStudy (anonymous):

2 4 5 8 25 50 62.50 100

OpenStudy (anonymous):

you still there @texaschic101?

OpenStudy (texaschic101):

First we find the slope by using the slope formula : slope(m) = (y2 - y1) / (x2 - x1) take 2 sets of points (2,25)...x1 = 2 and y1 = 25 (4,50)...x2 = 4 and y2 = 50 now we sub into our formula slope(m) = (50 - 25) / (4 - 2) slope(m) = 25/2 which reduces to 12.5 now we just use y = mx + b and since we found m, we must now find b, the y intercept pick a set of points...(2,25) ...x = 2 and y = 25 slope(m) = 12.5 now we sub 25 = 12.5(2) + b 25 = 25 + b -- subtract 25 from both sides 25 - 25 = b 0 = b so your equation is : y = 12.5x + 0....which is the same as y = 12.5x ok...I am not getting any of your answer choices. Is the problem written correctly ?

OpenStudy (anonymous):

100% sure :/

OpenStudy (texaschic101):

let me try it again...I will use different points...give me a sec

OpenStudy (anonymous):

ok

OpenStudy (texaschic101):

slope(m) = (y2 - y1) / (x2 - x1) (4,50)...x1 = 4 and y1 = 50 (8,100)...x2 = 8 and y2 = 100 now sub slope(m) = (100 - 50) / (8 - 4) slope(m) = 50/4 which equals 12.5 nope..I am gonna get the same answer...y = 12.5x sorry

OpenStudy (anonymous):

well thats confusing :/

OpenStudy (texaschic101):

you sure the answer choices are written correctly ??

OpenStudy (anonymous):

thanks anyway

OpenStudy (texaschic101):

sorry..I do not see what I did wrong :(

OpenStudy (anonymous):

A. Y= 11.5x B. Y=12.25x C. Y=12.75x Y=13x those are it, I can choose more than 1

OpenStudy (texaschic101):

I do not see how it can be anything else but y = 12.5x.....we better get someone else over here to look

OpenStudy (texaschic101):

@hartnn ..can you check this please

OpenStudy (texaschic101):

lets check it... (2,25)...x = 2 and y = 25 now sub into y = 12.5x 25 = 12.5(2) 25 = 25 (correct) I believe your answer choices are wrong in your book

hartnn (hartnn):

The table represents Relationship B.

OpenStudy (igreen):

None of them are right..

hartnn (hartnn):

y =12.5x is for B and A has greater rate

hartnn (hartnn):

so y=12.75x and y=13x both could represent A :)

OpenStudy (igreen):

Ohhh that's right!!

OpenStudy (igreen):

The answer is gonna be the answer with a greater slope than 12.5..

OpenStudy (igreen):

Which one's greater than 12.5? @Samus_Aran

OpenStudy (igreen):

Which one's are* greater than 12.5?

OpenStudy (igreen):

@hartnn I thought we weren't supposed to give answers out..

OpenStudy (anonymous):

thank you @hartnn

OpenStudy (anonymous):

D, C, and B

hartnn (hartnn):

not B